# Tropical Gradient Descent

### v1

@@ 2.2.3 Tropical Location Problems
The $\triangle$-star-convex functions defined above are generally a measure of closeness to the kernel $\mathbf{v}$, but in minimising a statistical loss function we look at closeness to a sample of points. A location problem (Laporte et al., (2019)) measures the closeness to a sample rather than a single kernel.
Definition 14 (Max(/min)-tropical location problem (Comăneci, (2023))).
Consider a dataset $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{n}/\mathbb{R}\mathbf{1}$. For $k\leq K$, let $h_{k}$ be a $\triangle_{\max}$-star-quasi-convex function with kernel $\mathbf{x}_{k}$, and let $g:\mathbb{R}^{m}\rightarrow\mathbb{R}$ be an increasing function. A max-tropical location problem is the minimisation of an objective function $f:\mathbb{R}^{n}/\mathbb{R}\mathbf{1}\rightarrow\mathbb{R}$ given by $h=g(h_{1},\dots,h_{m})$. A min-tropical location problem is defined similarly, where the $h_{k}$ are $\triangle_{\min}$-star-quasi-convex.
The main result by Comăneci, (2023) is the following; tropical location problems contain a minimum in the tropical convex hull of its dataset $X$.
Theorem 15 (Theorem 20, Comăneci, (2023)).
Let $f$ be the objective function of a max-tropical location problem. Then there is a minimum of $f$ belonging to $\operatorname{tconv}_{\min}(X)$. Conversely, min-tropical location problems have minima belonging to $\operatorname{tconv}_{\max}(X)$.
We can now revisit our motivating example (Definition 8), showing that tropical linear regression is not just a min-tropical location problem, but also demonstrates tropical quasi-convexity.
@@
The tropical linear regression problem is a tropically quasi-convex $\min$-tropical location problem.
Proof.
@@
These are min-tropical location problems with:
@@
This $g$ and $\gamma$ are increasing and each $h_{k}$ has kernel $x_{k}$, satisfying the conditions for a min-tropical location problem.
As the $h_{k}$ are $\triangle_{\min}$-star-convex, their sublevel sets are min-tropically convex. This is preserved when taking a maximum over samples, so $f$ also has $\min$-tropically convex sublevel sets. ∎
As mentioned in Lemma 13, $\triangle_{\min}$-star-quasi-convexity implies tropical quasi-convexity. However, the tropical linear regression problem is a case in which the reverse implication does not hold; for any sample of size $\geq 1$, the linear regression loss function is not $\triangle_{\min}$-star-quasi-convex.
@@ 2.3 Tropical Data Science
We now review various tropical statistical optimisation problems which have been introduced in the literature in recent years, and which are used to test our tropical descent methodology. We note the varying degrees of convexity demonstrated by each of them.
These statistical concepts and methods showcase the breadth of applicability of our proposed methodology: optimisation is a central task in computing important statistical quantities, tasks in both descriptive and inferential statistics, and distances between probability measures. Adapting these statistical tasks to the setting of tropical geometry is not straightforward and it is not immediate that a single tropical optimisation method can be used to find solutions to the optimisation problems.
@@ 2.3.1 Centrality Statistics
Our first problems of interest, Fermat–Weber Points and Fréchet means, are centrality statistics for data on a general metric space; Fermat–Weber points (Wesolowsky, (1993)) act as a generalised median, while Fréchet means (Fréchet, (1948)) are a generalisation of the mean.
Definition 17 (Tropical Fermat–Weber Points (Lin & Yoshida, (2018))).
Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ be our dataset. A solution to the following optimisation problem is a tropical Fermat–Weber point:
@@
Definition 18 (Tropical Fréchet means (Jazlan et al., (2024+); Monod et al., (2022))).
Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ be our dataset. A solution to the following optimisation problem is a tropical Fréchet mean:
@@
The Fermat–Weber and Fréchet mean problems are both max-tropical and min-tropical location problems, as the tropical metric $d_{\mathrm{tr}}$ is both $\triangle_{\min}$ and $\triangle_{\max}$-star-quasiconvex. In fact, they are both classically convex and hence act as a benchmark against which we will test our gradient methods.
@@ 2.3.2 Wasserstein Projections
The next statistical optimisation problem was motivated by Cai & Lim, (2022), and looks to identify a tropical projection which minimises the Wasserstein distance between samples in different tropical projective torii.
Definition 19 (Tropical Wasserstein Projections (Talbut et al., (2023))).
Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1},Y=\{\mathbf{y}_{1},\dots,\mathbf{y}_{K}\}\subset\mathbb{R}^{M}/\mathbb{R}\mathbf{1}$ be our datasets. Let $J_{1},\dots,J_{M}$ be a non-empty partition of $[N]$. Then the tropical $p$-Wasserstein projection problem is the minimisation of the following objective function:
$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f_{p}(\mathbf{t})=\left(\frac{1}{K}\sum_{k\leq K}\|[\max_{i\in J_{j}}(\mathbf{x}_{k}+\mathbf{t})_{i}]_{j}-\mathbf{y}_{k}\|_{\mathrm{tr}}^{p}\right)^{1/p}$
The tropical $\infty$-Wasserstein projection problem is the minimisation of:
$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f_{\infty}(\mathbf{t})=\max_{k\leq K}\|[\max_{i\in J_{j}}(\mathbf{x}_{k}+\mathbf{t})_{i}]_{j}-\mathbf{y}_{k}\|_{\mathrm{tr}}$
@@
The tropical $p,\infty$-Wasserstein projections are max-tropical location problems, and the $\infty$-Wasserstein projection has $\max$-tropically convex sublevel sets.
Proof.
Let $\mathbf{z}_{k}=(x_{ki}-y_{kj})_{i\leq N}$ where $j$ is such that $i\in J_{j}$. We then have that
$\|[\max_{i\in J_{j}}(\mathbf{x}_{k}+\mathbf{t})_{i}]_{j}-\mathbf{y}_{k}\|_{\mathrm{tr}}$ $=\max_{l}[t_{l}+z_{kl}]+\max_{j}[\min_{i\in J_{j}}-t_{i}-z_{ki}]$
$=\max_{j}[\min_{i\in J_{j}}[\max_{l}[t_{l}+z_{kl}]-t_{i}-z_{ki}]]$
@@
$f_{p}(\mathbf{t})$ $=\left(\frac{1}{K}\sum_{k\leq K}\left(\max_{j}[\min_{i\in J_{j}}[\max_{l}[t_{l}+z_{kl}]-t_{i}-z_{ki}]]\right)^{p}\right)^{1/p}$
$f_{\infty}(\mathbf{t})$ $=\max_{k\leq K}\max_{j}[\min_{i\in J_{j}}[\max_{l}[t_{l}+z_{kl}]-t_{i}-z_{ki}]]$
@@
$h_{k}(\mathbf{t})$ $=\hat{\gamma}(\mathbf{t}+\mathbf{z}_{k}),$
@@
The $g_{p},g_{\infty}$ and $\gamma$ are increasing and each $h_{k}$ has kernel $-z_{k}$, satisfying the conditions for a max-tropical location problem.
As the $h_{k}$ are $\triangle_{\max}$-star-convex, their sublevel sets are max-tropically convex. This is preserved when taking a maximum over samples, so $f_{\infty}$ also has $\max$-tropically convex sublevel sets. ∎
@@ 3 Steepest Descent with the Tropical Norm
We treat optimisation over $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ as optimisation over $\mathbb{R}^{N}$ with the knowledge that $f$ is constant along $\mathbf{1}$-rays. We assume that for all $\mathbf{t}$, we can compute some gradient $\nabla f(\mathbf{t})$ which is contained in the generalised Clarke derivative of $\mathbf{t}$.
Definition 21 (Generalised Clarke Derivative (2.13, I.I of Kiwiel, (2006))).
@@
Then the generalised Clarke derivative is given by
@@ 3.1 Proposed Method
In defining tropical descent, we refer back to the motivation behind classical gradient descent; that is, the steepest descent direction.
Definition 22 (Steepest Descent Direction (Boyd & Vandenberghe, (2004))).
@@
$\mathbf{d}=\operatorname{argmax}\{\nabla f(\mathbf{t})^{\top}\mathbf{v}:\|\mathbf{v}\|=1\}.$
With respect to the Euclidean norm, the normalised gradient is a steepest descent direction. When we use a the tropical norm, the steepest descent directions are characterised by the following lemma.
@@
Suppose $\nabla f\neq 0$. Then direction $\mathbf{v}$ is a normalised tropical steepest descent direction iff
@@
Proof.
WLOG, we assume $0\leq v_{i}\leq 1$. Then
$\nabla f^{\top}\mathbf{v}\leq\sum_{\nabla f_{i}>0}\nabla f_{i}$
This is achieved if and only if $v_{i}=1$ for all $i$ such that $\nabla f_{i}>0$, and $v_{i}=0$ for all $i$ such that $\nabla f_{i}<0$. We note that any such $v_{i}$ satisfies $\|\mathbf{v}\|_{{\rm tr}}=1$ as $\sum\nabla f_{i}=0,\nabla f_{i}\neq 0$, and hence $\nabla f_{i}$ has positive and negative coordinates. ∎
This lemma uniquely defines $\mathbf{d}$ if and only if, for every coordinate $i$, $\nabla f_{i}\neq 0$. To uniquely define a descent direction for any $\mathbf{x}$, we take the direction $\mathbf{d}$ given by
$d_{i}=\begin{cases}1&\text{ if }\nabla f_{i}>0,\\ 0&\text{ otherwise.}\end{cases}$
We refer to this as a positive descent direction. In contrast,
$d_{i}=\begin{cases}-1&\text{ if }\nabla f_{i}<0,\\ 0&\text{ otherwise.}\end{cases}$
defines a negative descent direction.
Throughout the rest of this paper, we use the positive descent direction as convention, and will specify when the negative descent direction is used instead. Tropical descent therefore refers to the method outlined by Algorithm 1, for some specified step size rule, while we refer to classical gradient descent as classical descent.
@@
     $\mathbf{g}_{n}\leftarrow\nabla f(\mathbf{t}_{n-1})$;
     $\mathbf{d}_{n}\leftarrow[\mathbf{g}_{n}>0]_{i\leq N}$;
     $\mathbf{t}_{n}\leftarrow\mathbf{t}_{n-1}-a_{n}\mathbf{d}_{n}$;

### v2

@@ 2.2.3 Tropical Location Problems
The $\triangle_{\min}$-star-convex functions defined above are generally a measure of closeness to the kernel $\mathbf{v}$, but in minimizing a statistical loss function we look at closeness to a sample of points. A location problem (Laporte et al.,, 2019) measures the closeness to a sample rather than a single kernel.
Definition 16 (Min(/max)-Tropical Location Problem (Comăneci,, 2024)).
Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ be our dataset. For $k\leq K$, let $h_{k}$ be a $\triangle_{\min}$-star-quasi-convex function with kernel $\mathbf{x}_{k}$, and let $g:\mathbb{R}^{K}\rightarrow\mathbb{R}$ be increasing in every coordinate. A min-tropical location problem is the minimization of an objective function $f:\mathbb{R}^{N}/\mathbb{R}\mathbf{1}\rightarrow\mathbb{R}$ given by $f=g(h_{1},\dots,h_{K})$. A max-tropical location problem is defined similarly, where the $h_{k}$ are $\triangle_{\max}$-star-quasi-convex.
The main result of (Comăneci,, 2024) is the following: tropical location problems contain a minimum in the tropical convex hull of its dataset $X$.
Theorem 17 (Theorem 20 of (Comăneci,, 2024)).
Let $f$ be the objective function of a min-tropical location problem. Then there is a minimum of $f$ belonging to $\operatorname{tconv}_{\max}(X)$. Conversely, max-tropical location problems have minima belonging to $\operatorname{tconv}_{\min}(X)$.
We can now revisit our motivating example (Definition 9), showing that tropical linear regression is not just a min-tropical location problem, but also demonstrates tropical quasi-convexity in that it has tropically convex sublevel sets.
@@
The tropical linear regression problem is a $\min$-tropical location problem with tropically convex sublevel sets.
Proof
@@
This is a min-tropical location problem with:
@@
This $g$ and $\gamma$ are increasing in each coordinate and each $h_{k}$ has kernel $\mathbf{x}_{k}$, satisfying the conditions for a min-tropical location problem.
As the $h_{k}$ are $\triangle_{\min}$-star-convex, their sublevel sets are min-tropically convex by Proposition 15. This is preserved when taking a maximum over samples, so $f$ also has $\min$-tropically convex sublevel sets. ∎
As mentioned in Proposition 15, $\triangle_{\min}$-star-quasi-convexity implies tropical quasi-convexity. However, the tropical linear regression problem is a case in which the reverse implication does not hold; in general, the linear regression loss function is not $\triangle_{\min}$-star-quasi-convex.
@@ 2.3 Tropical Data Science
While young, the field of tropical data science is motivated by the crucial question of statistical learning from the increasing volume of phylogenetic tree data (Kahn,, 2011). In contrast to the BHV (Billera et al.,, 2001) or Robinson–Foulds metrics (Robinson & Foulds,, 1981), the tropical interpretation of tree space provides both interpretable geometry and computational efficiency (Steel & Penny,, 1993; Monod et al.,, 2018). Many of the statistical methods defined for the tropical setting are framed as optimisation problems (Barnhill et al.,, 2024b), and in this subsection we review some such tropical statistical optimization problems which we use to test our tropical descent methodology. We note the varying degrees of convexity demonstrated by each of them. While we restrict our considerations through the rest of this paper to statistics which arise from tropical location problems, in Appendix A we discuss other tropical statistical optimization problems.
@@ 2.3.1 Centrality Statistics
Our first problems of interest, Fermat–Weber points and Fréchet means, are centrality statistics for data on a general metric space; Fermat–Weber points (Drezner et al.,, 2004) act as a generalized median, while Fréchet means (Fréchet,, 1948) are a generalization of the mean.
Definition 19 (Tropical Fermat–Weber Points (Lin & Yoshida,, 2018)).
Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ be our dataset. A solution to the following optimization problem is a tropical Fermat–Weber point:
@@
Definition 20 (Tropical Fréchet means (Monod et al.,, 2018)).
Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ be our dataset. A solution to the following optimization problem is a tropical Fréchet mean:
@@
The Fermat–Weber and Fréchet mean problems are both max-tropical and min-tropical location problems, as the tropical metric $d_{\mathrm{tr}}$ is both $\triangle_{\min}$ and $\triangle_{\max}$-star-quasi-convex. In fact, they are also both classically convex and hence act as a benchmark against which we can test our gradient methods.
@@ 2.3.2 Wasserstein Projections
The next statistical optimization problem was motivated by Cai & Lim, (2022), and looks to identify a tropical projection which minimizes the Wasserstein distance between samples in different tropical projective tori.
Definition 21 (Tropical Wasserstein Projections (Talbut et al.,, 2023)).
Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ be a dataset in one space, and $Y=\{\mathbf{y}_{1},\dots,\mathbf{y}_{K}\}\subset\mathbb{R}^{M}/\mathbb{R}\mathbf{1}$ be a dataset in a different space. Let $J_{1},\dots,J_{M}$ be a non-empty partition of $[N]$. Then the tropical $p$-Wasserstein projection problem is the minimization of the following objective function:
$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f_{p}(\mathbf{t})=\left(\frac{1}{K}\sum_{k\leq K}\|[\max_{i\in J_{j}}(\mathbf{x}_{k}-\mathbf{t})_{i}]_{j}-\mathbf{y}_{k}\|_{\mathrm{tr}}^{p}\right)^{1/p}$
The tropical $\infty$-Wasserstein projection problem is the minimization of:
$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f_{\infty}(\mathbf{t})=\max_{k\leq K}\|[\max_{i\in J_{j}}(\mathbf{x}_{k}-\mathbf{t})_{i}]_{j}-\mathbf{y}_{k}\|_{\mathrm{tr}}$
@@
The tropical $p,\infty$-Wasserstein projections are min-tropical location problems, and the $\infty$-Wasserstein projection has $\min$-tropically convex sublevel sets.
Proof Let $\mathbf{z}_{k}=(x_{ki}-y_{kj})_{i\leq N}$ where $j$ is such that $J_{j}$ is the unique partition set containing $i$. We then have that
$\|[\max_{i\in J_{j}}(\mathbf{x}_{k}-\mathbf{t})_{i}]_{j}-\mathbf{y}_{j}\|_{\mathrm{tr}}$ $=\max_{b}\max_{a\in J_{b}}[z_{ka}-t_{a}]-\min_{j}\max_{i\in J_{j}}[z_{ki}-t_{i}]$
$=\max_{a}[z_{ka}-t_{a}]+\max_{j}\min_{i\in J_{j}}[t_{i}-z_{ki}]$
$=\max_{j}[\min_{i\in J_{j}}[t_{i}-z_{ki}-\min_{a}[t_{a}-z_{ka}]]]$
@@
$f_{p}(\mathbf{t})$ $=\left(\frac{1}{K}\sum_{k\leq K}\left(\max_{j}[\min_{i\in J_{j}}[t_{i}-z_{ki}-\min_{a}[t_{a}-z_{ka}]]]\right)^{p}\right)^{1/p}$
$f_{\infty}(\mathbf{t})$ $=\max_{k\leq K}\max_{j}[\min_{i\in J_{j}}[t_{i}-z_{ki}-\min_{a}[t_{a}-z_{ka}]]]$
@@
$h_{k}(\mathbf{t})$ $=\hat{\gamma}(\mathbf{t}-\mathbf{z}_{k}),$
@@
The $g_{p},g_{\infty}$ and $\gamma$ are increasing in each coordinate and each $h_{k}$ has kernel $\mathbf{z}_{k}$, satisfying the conditions for a min-tropical location problem.
As the $h_{k}$ are $\triangle_{\min}$-star-convex, their sublevel sets are min-tropically convex by Proposition 15. This is preserved when taking a maximum over samples, so $f_{\infty}$ also has $\min$-tropically convex sublevel sets. ∎
@@
Figure 4 shows a heat map for the $\infty$-Wasserstein projection objectives. As in the tropical linear regression problem, it is piecewise linear, non-convex and has sparse gradient.
@@ 3 Steepest Descent with the Tropical Norm
We treat optimization over $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ as optimization over $\mathbb{R}^{N}$ with the knowledge that $f$ is constant along $\mathbf{1}$-rays. We assume we can compute some gradient-like function $\nabla f:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ which is satisfying Assumption 23.
Assumption 23.
Let $f=g(h_{1},\dots,h_{K})$ be a min-tropical location problem. Then we have that
$\nabla f(\mathbf{t})_{j}<0\Rightarrow\exists\,k\text{ s.t. }j\in\text{argmin}_{i}(t_{i}-x_{ki}).$
This is a very general assumption, which holds for some of the most natural gradient-like functions. For example, if $\nabla f(\mathbf{t})$ is given by some Clarke gradient, that is, some element of the generalized Clarke derivative of $f$, then Lemma 25 shows that Assumption 23 holds.
Definition 24 (Generalized Clarke Derivative (2.13, I.I of Kiwiel, (1985))).
@@
Then the generalized Clarke derivative is given by
@@
Lemma 25.
Suppose $f=g(h_{1},\dots,h_{K})$ is min-tropical location problem, and for all $\mathbf{t}$, $\nabla f(\mathbf{t})$ is some element of $\partial f(\mathbf{t})$. Then Assumption 23 holds.
Proof Suppose that $\nabla f(\mathbf{t})_{j}<0$. Then there is a sequence of differentiable $\mathbf{y}_{\ell}$ satisfying $\nabla f(\mathbf{y}_{\ell})_{j}<0$, $\mathbf{y}_{\ell}\rightarrow\mathbf{t}$. Hence by the chain rule, as $g$ is increasing in every $h_{k}$, there is some $h_{k}$ which is strictly decreasing in the $j^{th}$ coordinate at $\mathbf{y}_{\ell}$. WLOG, the same $h_{k}$ is strictly decreasing in the $j^{th}$ coordinate for all $\mathbf{y}_{\ell}$. As $\gamma$ is increasing in every coordinate, $h_{k}=\hat{\gamma}(\mathbf{y}_{\ell}-\mathbf{x}_{k})$ is strictly decreasing in the $j^{th}$ coordinate if and only if $j\in\text{argmin}_{i}(y_{\ell i}-x_{ki})$. This is preserved through the limit as $\ell\rightarrow\infty$, so we conclude that $j\in\text{argmin}_{i}[t_{i}-x_{i}]$. ∎
In practice, rather than computing elements of the generalized Clarke derivative, we use PyTorch automatic differentiation (Paszke et al.,, 2019). The derivatives computed by PyTorch are dependent on the computational representation of a function, but are designed to respect the chain rule. We must therefore ensure that our computational representation of $g$ has non-negative derivatives in each $h_{k}$, while each $h_{k}$ is computed such that its $t_{j}$ derivative is negative only if $i\in\text{argmin}_{i}(t_{i}-x_{ki})$. For such a computational representation of $f$, Assumption 23 will hold by the chain rule. For all the tropical location problems we consider in this paper, we find their most natural computational representation gives rise to derivatives $\nabla f$ satisfying Assumption 23.
@@ 3.1 Proposed Method
In defining tropical descent, we refer back to the motivation behind classical gradient descent—steepest descent directions.
Definition 26 (Steepest Descent Direction (Boyd & Vandenberghe,, 2004)).
@@
$\mathbf{d}=\text{argmin}\{\nabla f(\mathbf{t})^{\top}\mathbf{v}:\|\mathbf{v}\|=1\}.$
With respect to the Euclidean norm, the normalized derivative is a steepest descent direction. When we use a the tropical norm, the steepest descent directions are characterized by the following lemma.
@@
Suppose $\nabla f\neq\mathbf{0}$. Then the direction $\mathbf{v}$ is a normalized tropical steepest descent direction iff
@@
Proof WLOG, we assume $0\leq v_{i}\leq 1$. Then
$\nabla f^{\top}\mathbf{v}\geq\sum_{i:\,\nabla f_{i}<0}\nabla f_{i}$
This is achieved if and only if $v_{i}=1$ for all $i$ such that $\nabla f_{i}<0$, and $v_{i}=0$ for all $i$ such that $\nabla f_{i}>0$. We note that any such $v_{i}$ satisfies $\|\mathbf{v}\|_{{\rm tr}}=1$ as $\sum_{i}\nabla f_{i}=0$ by the equivalence relation of the tropical projective torus, and $\nabla f_{i}\neq\mathbf{0}$ so $\nabla f$ has positive and negative coordinates. ∎
Lemma 27 uniquely defines a descent direction $\mathbf{d}$ if and only if, for every coordinate $i$, $\nabla f_{i}\neq 0$. To uniquely define a descent direction for any $\mathbf{t}$, we take the direction $\mathbf{d}$ given by
$d_{i}=\begin{cases}1&\text{ if }\nabla f_{i}<0,\\ 0&\text{ otherwise.}\end{cases}$
We refer to this as a min-tropical descent direction. In contrast,
$d_{i}=\begin{cases}-1&\text{ if }\nabla f_{i}>0,\\ 0&\text{ otherwise.}\end{cases}$
defines a max-tropical descent direction.
Throughout the rest of this paper, we use the min-tropical descent direction as convention, and will specify when the max-tropical descent direction is used instead. Tropical descent therefore refers to the method outlined by Algorithm 1, for some specified step size sequence $(a_{n})_{n\geq 1}$, while we refer to classical gradient descent as classical descent.
@@
     $\mathbf{g}_{ni}\leftarrow\nabla f_{i}(\mathbf{t}_{n-1})$;
     $\mathbf{d}_{ni}\leftarrow\mathbf{1}_{\mathbf{g}_{ni}<0}$;
     $\mathbf{t}_{n}\leftarrow\mathbf{t}_{n-1}+a_{n}\mathbf{d}_{n}$;
We note that over a large number of steps, Algorithm 1 can produce large $t_{ni}$ values as each coordinate is non-decreasing in $n$. If this results in overflow or precision errors, a normalisation operation can be included in each step, such as subtracting $\frac{1}{N}\sum_{i}t_{ni}$ from each $t_{ni}$. The function $f$ is invariant under such a translation due to the equivalence relation on the tropical projective torus.
In (Barnhill et al.,, 2024a, b), classical descent is implemented for the computation of Fermat–Weber points; this is the most natural choice as the Fermat–Weber problem is classically convex. However, the linear regression and Wasserstein projection problems lack classical convexity, and hence do not consistently produce accurate solutions via classical descent. It is these tropically convex problems for which tropical descent is necessary.